1. The IMF forecasted Australia’s real GDP at $A1,730 billion in 2017 and $A1,782 billion in 2018, and Australia’s population at 24.6 million in 2017 and 25.0 million in 2018. Calculate:
  2. The growth rate of real GDP
  3. The growth rate of real GDP per person
  4. The approximate number of years it will take for real GDP per person in Australia to double if the current growth rate of real GDP is maintained.
1
Expert's answer
2022-02-23T13:02:28-0500

(a).Growth rate of real GDP="(\frac{Real GDP_{year2}}{RealGDP_{base year}})" -1

="(\frac{Real GDP_{2018}}{Real GDP_{2017}})" -1

="\frac{1782}{1730}"-1=1.030-1

=0.03

=3%

therefore growth rate is 3%.

(b). growth rate per person

Real GDP per capita (2017)="\frac{Real GDP}{Population}"

="\frac {1730billion}{24.6million}" = $70325.20 billion

Real GDP per capita(2018)="\frac{Real GDP}{Population}"

="\frac{1782billion}{25million}" ="\frac{1782billion}{0.025billion}"

=$71280 billion

so growth rate will be "\frac{70325.20-71280}{70325.20}" "\times{100}"

"\frac{-954.8}{70325.20}\times{100}"

=-0.0136"\times{100}"

-1.36% thus growth rate of real person is -1.36%

(c).period to double the current real GDP growth rate is; we us ethe rule of 70, therefore;

"\frac{70}{1.36}" =51.47 years.

thus it took approximately 51.47 years to double the real GDP per person